# Beyond Diagonal RIS: Passive Maximum Ratio Transmission and Interference Nulling Enabler

### v1

@@ V-C Simulation Results
Fig. 3 illustrates the received SINR at stage 1 of a BD-RIS with $K=5$, comparing relaxed and symmetric unitary solutions for BD-RIS MRT and BD-RIS Max-F. Specifically, the figure breakdowns (5) into the sum of signal powers, sum of interference and Frobenius norm of the equivalent channel for $N=64$ in a stacked bar chart. Also, it depicts the power ratio of the symmetric unitary solution to the relaxed solution. It is seen that the relaxed BD-RIS Max-F achieves the highest Frobenius norm having the highest bar. Nonetheless, the symmetric unitary solution for BD-RIS Max-F does not guarantee the maximization of the Frobenius norm as seen in the figure, where the symmetric unitary BD-RIS MRT is better. In addition, the sum of signal powers for the relaxed and symmetric unitary BD-RIS MRT are higher than that of BD-RIS Max-F as expected, whereas the sum of interference is higher for BD-RIS Max-F compared to BD-RIS MRT as the latter only maximizes the diagonal entries of the equivalent channel but the former maximizes all entries. Such observation indicates that BD-RIS MRT would have a better SINR at the users compared to BD-RIS Max-F. Furthermore, although the symmetric unitary projection degrades the sum of signal powers for both BD-RIS Max-F and BD-RIS MRT, it reduces the sum of interference for BD-RIS MRT at a higher rate such that the SINR can be effectively improved, unlike BD-RIS Max-F which reduces both the sum of signal powers and sum of interference at the same rate.
Figure 4: Sum rate vs. transmit power for BD-RIS and D-RIS considering Null with random initialization, where $K=5$ and $(2K-1)\leq N\leq 2^{K+1}$.
(a) $N=31$
(b) $N=64$
Figure 5: Sum rate vs. transmit power for BD-RIS considering Null and MRT, where $K=5$ and a) $N=31$, b) $N=64$.
Fig. 4 compares the sum rate performance of AO to achieve interference nulling using BD-RIS and D-RIS [10]. Random initialization and uniform power allocation at the BS is assumed for both, where $K=5$ and $(2K-1)\leq N\leq 2^{K+1}$. It is shown that BD-RIS achieves better sum rate performance with a huge gap when $N=2K-1$, which is due to the fact that D-RIS needs $N>2K(K-1)$ to effectively achieve interference nulling with high probability. As $N$ grows towards $2^{K+1}$, the performance gap reduces accordingly, i.e., the performance gain of BD-RIS vanishes for $N>2K(K-1)$, which is justified as the interference is nulled for both BD-RIS and D-RIS considering the random initialization. This indicates other initialization points need to be investigated to show the BD-RIS performance gain over D-RIS. Furthermore, the snippet shows that for a given number of REs BD-RIS achieves an increasing sum rate by increasing $K$, unlike the D-RIS which achieves a reduced sum rate before saturation. This indicates that BD-RIS is a better candidate for interference nulling in dense MU-MISO due to its flexibility in designing $\boldsymbol{\Theta}$.
@@
Figure 6: Sum rate for the benchmarks and proposed BD-RIS designs with $P_{\max}=5$ dBm. a) vs. number of users, $N=5K$. b) vs. number of REs, $K=2$.
@@
Fig. 5 shows the sum rate of BD-RIS vs. the transmit power at the BS. It considers the proposed BD-RIS designs. Specifically, MRT and Null with MRT initialization (14) are considered. Also, the impact of different BS processes are considered. The figure considers $K=5$ users and $N\in\{31,64\}$ REs. While focusing on UP, Fig. 5 shows that MRT enhances the sum rate at low SNRs, a regime in which the performance is dominated by noise. This is due to its better channel alignment compared to Null. Nonetheless, the performance of MRT is saturated at moderate and high SNRs due to being interference-limited, unlike Null. Such saturation is observed for smaller SNRs when the number of REs is increased since increasing the number of REs improves the SNR in general. On the other hand, RM power allocation schemes improve the sum rate in general. Moreover, RM improves the sum rate of Null slightly at low SNRs and small BD-RIS, whereas its improvement for high SNRs and larg BD-RIS is negligible. Finally, MRT with ZF precoding at the BS outperforms Null with RM at the BS by $\sim 3$ dB at $20$ bps/Hz. Hence, MRT with ZF at the BS maintains low computational complexity while achieving a good sum rate performance at low and moderate SNRs and the best sum rate performance at high SNRs.
@@
Fig. 6 illustrates the sum rate for the proposed BD-RIS and the benchmarks, where $P_{\max}=5$ dBm. In Fig. 6a, the sum rate is illustrated vs. the number of users, where $N=5K$ REs. It can be seen that BD-RIS Joint serves as an upper bound for the sum rate for $N>2$, where the gap increases as the number of users increase. This is justified because it optimizes both the scattering matrix of the BD-RIS and the precoding matrix at the BS simultaneously to maximize the sum rate unlike other benchmarks and proposed solutions. For instance, at $K=10$, BD-RIS Joint achieves a sum rate of $14$ bps/Hz, whereas BD-RIS MRT with ZF at the BS achieves a sum rate of $10.3$ bps/Hz [^1] It is worth noting that the unitary constraint is only considered for BD-RIS Joint while the rest BD-RIS designs consider both symmetric and unitary constraints.. Moreover, it is noted that nMIMO curves have the same slope due to its limited power budget. Furthermore, nMIMO ZF achieves better sum rate than nMIMO MRT for such SNR. This is consistent with literature since the conventional ZF outperforms MRT at higher SNRs because ZF has the ability to minimize inter-user interference with slight power loss that is insignificant at such SNRs. Also, both proposed BD-RIS designs outperform BD-RIS Max-F and BD-RIS Max-l2 designs because the proposed BD-RIS designs have better interference management in multiuser systems. In addition, BS ZF are both performing poorly since BD-RIS is not utilized to enhance the poor channel gains stemming from the double-fading effect. Therefore, relying solely on the BS for precoding becomes inefficient. For instance, ZF wastes most of the power for channel inversion leaving very small power gains. Moreover, the performance of the proposed designs improve as the number of users/REs increase which is justified as the scattering matrix Frobenius norm increases as well indicating their effectiveness in handling dense multiuser scenarios. While the gap between BD-RIS MRT and BD-RIS Null with MRT initialization increases as the number of users increase, the latter can be a good candidate to achieve a reasonable sum rate performance meanwhile maintaining a negligible complexity at the BS. Furthermore, Fig. 6b shows the sum rate vs. the number of REs for $K=2$. It is seen that nMIMO curves achieve the best sum rate in this regime, while the specular reflection BD-RIS curves are both lower bounds and the justification holds from Fig. 6a. Furthermore, the gap between BD-RIS Joint and BD-RIS MRT is reduced significantly and it vanishes as the number of REs grows. This indicates that passive MRT with ZF precoding at the BS achieves local optimality while maintaining low complexity, and is considered a good candidate for BD-RIS with many REs. Also, the significance of BD-RIS Max-F and BD-RIS Max-l2 is shown for high number of REs with an advantage for BD-RIS Max-F. Their poor performance with relatively small $N$ is justified by the fact that the equivalent channel is ill-conditioned/poorly conditioned for such $N$ values and it improves as $N$ increases.
@@
Fig. 7 shows the time complexity considering the proposed BD-RIS and the benchmarks, where $K=2$ and $P_{\max}=5$ dBm. The worst-case scenario is considered for BD-RIS Joint and BD-RIS Null with MRT initialization, such that $t_{1}=50$, $t_{2}=20$, and $t=100$. As presented in Sec. III and Sec. V-B, BD-RIS MRT has the lowest time complexity while the BD-RIS Max-F has the highest. In addition, BD-RIS Max-l2 has a slightly higher time complexity than BD-RIS MRT because the former requires three SVD operations while the latter requires only one. Although BD-RIS Joint has a cubic time complexity, the complexity from the iterative algorithm is not negligible.
Figure 7: Time complexity for the benchmarks and proposed BD-RIS designs with $P_{\max}=5$ dBm vs. number of REs, $K=2$.
@@ VI Conclusions and Future Work
To conclude, this work proposed a passive multiuser beamforming for BD-RIS-enabled MU-MISO systems. The passive multiuser beamforing designs are inspired by MRT and interference nulling. The former passive beamforming is presented in closed-form while the latter is presented in an AO framework. At the BS side, optimized/uniform power allocation and active precoding are considered. It is seen that BD-RIS reduces the number of REs required to achieve interference nulling and achieves faster convergence compared to D-RIS. In addition, the passive MRT at BD-RIS with ZF precoding at the BS presents an effective low-complexity solution that simplifies channel estimation and achieves local optimality for MU-MISO scenarios with many REs.
For the future work, we will consider deriving a closed-form unitary symmetric scattering matrix to achieve interference nulling for the BD-RIS. In addition, we will analyze the average SINR and outage probability for the passive MRT. An extension to multi-antenna users will also be considered.

### v2

@@ V-C Simulation Results
Fig. 3 illustrates the received SINR at stage 1 of a fully-connected and group-connected BD-RISs with $K=5$, comparing relaxed and symmetric unitary solutions for MRT and Max-F. Specifically, the figure breakdowns (11) into the sum of desired signals’ powers and sum of interference signals’ powers for $N=16$ in a stacked bar chart. It is seen that the relaxed Max-F achieves the highest Frobenius norm having the highest bar. Nonetheless, the symmetric unitary solution for Max-F does not guarantee the maximization of the Frobenius norm as seen for the fully-connected case, where the symmetric unitary MRT is better. In addition, the sum of desired signals’ powers for the relaxed and symmetric unitary MRT are higher than that of Max-F as expected, whereas the sum of interference signals’ powers is higher for Max-F compared to MRT as the latter only maximizes the diagonal entries of the equivalent channel but the former maximizes all entries. Such observation indicates that MRT would have a better SINR at the users compared to Max-F. Furthermore, although the SVD-based symmetric unitary projection degrades the sum of desired signals’ powers for both Max-F and MRT, it reduces the sum of interference signals’ powers for MRT at a higher rate such that the SINR can be effectively improved, unlike Max-F which reduces both the sum of signal powers and sum of interference at the same rate.
@@
Figure 6: Sum rate for the benchmarks and proposed designs for fully-connected, group-connected and single-connected BD-RISs with $P_{\max}=5$ dBm. a) x-axis is number of users, $N=2K(K-1)$. b) x-axis is number of REs, $K=2$.
Fig. 4 compares the sum rate performance of interference nulling using single/group/fully-connected BD-RISs. The system uses MRT initialization and uniform power allocation at the BS, with $K=5$ and $N\in\{32,64\}$ [^1] Note that uniform power allocation is not optimal to maximize the sum rate. Nonetheless, it is considered in this figure to assess the sum rate performance of passive interference nulling using BD-RIS.. As shown in Fig. 4a, the fully-connected BD-RIS outperforms the single-connected BD-RIS for both $N=32$ and $N=64$ with a huge gap. This performance gap is linked to the fully-connected BD-RIS requiring fewer REs to effectively null the interference with high probability. Specifically, (33) shows that the fully-connected BD-RIS requires $N\geq 2K-1$, while the single-connected BD-RIS needs $N\geq 2K(K-1)$. Furthermore, Fig. 4b shows that increasing the number of users, while fixing the number of REs, does not necessarily improve the sum rate. For instance, when $N$ is sufficiently large, the fully-connected BD-RIS has an improved sum rate as more users are served. This occurs because, when $N\gg 2K-1$, interference is nulled successfully using the required $12K-$ 1 REs, leaving a substantial potion of the remaining REs to enhance the signal power. In contrast, both the group-connected and single-connected BD-RISs show a degraded sum rate as the number of users increases. This degradation arises because $N$ is not sufficiently large to achieve $N\gg 2K(K-1)$ and $N\gg\left\lceil\frac{4K(K-1)}{1+N_{g}}\right\rceil$. Consequently, most of the REs are used to null the interference, i.e., $2K(K-1)$ and $\left\lceil\frac{4K(K-1)}{1+N_{g}}\right\rceil$ REs, leaving only a small portion of REs to enhance the signal power.
@@
Fig. 5 shows the sum rate of single/group/fully-connected BD-RISs as a function of transmit power at the BS. It considers the proposed BD-RIS designs, specifically considering the MRT and Null schemes. Additionally, the impact of different stage 2 processes at the BS is considered. The highlighted scenario in the figure is $K=5$ users and $N=64$ REs. Focusing on UP, Fig. 5 shows that MRT outperforms Null at low SNRs, where performance performance is primarily dominated by noise. This is because MRT provides a better channel alignment compared to Null. However, at moderate and high SNRs, MRT becomes interference-limited leading its performance to saturate, unlike Null, which inherently eliminates the interference. On the other hand, RM power allocation schemes improve the sum rate of MRT in general. While RM also enhances the sum rate of Null at low SNRs, the improvement at high SNRs is negligible. Finally, considering the fully-connected BD-RIS as an example, MRT with ZF precoding at the BS outperforms Null with RM at the BS by $1.3$ dB at $20$ bps/Hz. Thus, MRT with ZF at the BS achieves a balance between low computational complexity and good sum rate performance at low and moderate SNRs, while offering the best sum rate performance at high SNRs.
@@
Fig. 6 presents the sum rate performance for both the proposed designs and the benchmarks across single/group/fully-connected BD-RISs, where $P_{\max}=5$ dBm. In Fig. 6a, the sum rate is illustrated as a function of number of users, where $N=2K(K-1)$ REs. It can be seen that Joint serves as an upper bound, with the gap between the upper bound and the proposed designs varying based on the BD-RIS architecture. For instance, the gap is extremely tight for the fully-connected BD-RIS since $N\gg 2K-1$, meaning that $N$ is sufficiently large compared to the bound in (33). In contrast, the group-connected and single-connected BD-RISs exhibit a much larger gap from the upper bound, as $N$ is not sufficiently large when compared to the bound in (33). For instance, at $K=8$, the fully-connected BD-RIS with Joint achieves a sum rate of $28.3$ bps/Hz, whereas MRT with ZF at the BS achieves a sum rate of $27.7$ bps/Hz [^2] It is worth noting that Joint only considers (9d) constraint, while the rest BD-RIS designs consider both (9c) and (9d) constraints.. Additionally, both proposed BD-RIS designs outperform Max-F designs because the proposed BD-RIS designs have better interference management in multiuser systems. In addition, Spec-ZF and Spec-MRT are both performing poorly, particularly for the fully-connected BD-RIS. This is because these designs fail to leverage the BD-RIS to mitigate the poor channel gains caused by the double-fading effect, making reliance on the BS solely for precoding inefficient. For instance, ZF wastes most of the power for channel inversion leaving very small power gains. Moreover, the performance of the proposed designs improve as the number of users/REs increase which is justified by the growing Frobenius norm of the scattering matrix, indicating their effectiveness in handling dense multiuser scenarios. While the gap between MRT and Null with MRT initialization increases as the number of users increase, the latter can be a good candidate to achieve a reasonable sum rate performance meanwhile maintaining a negligible complexity at the BS.
Furthermore, Fig. 6b shows the sum rate as a function of number of REs for $K=2$. It is seen that Joint serves as an upper bound here too, where the gap between the upper bound and the proposed designs is very tight given that the range of $N$ is sufficiently large for all BD-RIS architectures given $K=2$. This indicates that passive MRT with ZF precoding at the BS achieves local optimality while maintaining low complexity, and is considered a good candidate for BD-RIS with many REs. The specular reflection BD-RIS curves on the other hand are both lower bounds and the justification holds from Fig. 6a. Also, the significance of Max-F is shown for high number of REs. Their poor performance with relatively small $N$ is justified by the fact that the equivalent channel is ill-conditioned/poorly conditioned for such $N$ values and it improves as $N$ increases.
@@
Fig. 7 shows the computational complexity considering the proposed designs and benchmarks for single/group/fully-connected, where $K=2$. The worst-case scenario is considered for Joint and Null with MRT initialization, such that $t_{1}=50$, $t_{2}=20$, $t_{3}=5$, and $t=100$. As presented in Table I and Sec. V-V-B, MRT has the lowest time complexity while the Max-F has the highest. Although Joint has a cubic time complexity, the complexity from the iterative algorithm is not negligible. Furthermore, Null has a lower complexity compared to both benchmarks.
Figure 7: Computational complexity as a function of number of REs. The proposed designs and benchmarks are considered for fully-connected “FC”, group-connected “GC” and single-connected “SC” BD-RISs, where $K=2$.
@@ VI Conclusions and Future Work
To conclude, this work proposed a passive multiuser beamforming for single/group/fully-connected BD-RIS-enabled MU-MISO systems. The passive multiuser beamforing designs are inspired by MRT and interference nulling. The former passive beamforming is presented in closed-form while the latter is presented in an AO framework. At the BS side, optimized/uniform power allocation and active precoding are considered. It is seen that the fully-connected BD-RIS reduces the number of REs required to achieve interference nulling and achieves faster convergence compared to D-RIS. In addition, the passive MRT at BD-RIS with ZF precoding at the BS presents an effective low-complexity solution that simplifies channel estimation and achieves local optimality for MU-MISO scenarios with many REs.
For the future work, we will consider deriving a closed-form unitary symmetric scattering matrix to achieve interference nulling for the BD-RIS. In addition, we will analyze the average SINR and outage probability for the passive MRT. An extension to multi-antenna users will also be considered. Furthermore, we will generalize the BD-RIS scattering matrix design and BS precoder design to account for the imperfect CSI.